Suppose the absolute values of the intercept and slope of the demand function



are approximated to be ten (10) and three (3) respectively. If the absolute



values of the intercept and slope of the supply function are assessed to be six (6),



and five (5) respectively, calculate equilibrium price and quantity (4

1
Expert's answer
2021-11-07T19:40:19-0500

A general demand function is written as :

"Qd=a-bP"

Qd-Quantity demanded

a-Intercept :It is Qd when P is 0

b-Slope : It measures how much Qd changes when P changes.

P-Price

Given:

"Qd=a-bP"

So, the demand function is

"Qd=10-3P"

 


A general supply function is written as :

"Qs=c+dP"

Qs-Quantity supplied

c-Intercept :It is Qs when P is 0

d-Slope : It measures how much Qs changes when P changes.

P-Price

Given:

"c=6\\

d=5"

So, the supply function is

"Qs=6+5P"


At equilibrium: "Qd=Qs"

"Qd=10-3P\\

Qs=6+5P\\

Qd=Qs\\

10-3P=6+5P\\

8P=4\\

P^*=0.5\\

Q^*=8.5"